# Hybrid inflation and gravitational waves from accidentally light scalars

### v1

@@ Appendix B A model with disconnected minima, giving rise to unstable domain walls
Let us first consider the case $\mu_{\phi}^{2}>0$ and $\mu_{\chi}^{2}<0$. The minimum is at $\langle\phi^{2}\rangle=v^{2}\equiv\mu_{\phi}^{2}/\lambda_{\phi}$, $\langle\chi\rangle=0$, where the mass spectrum resembles the one of the minimal model of section 2. After the spontaneous breaking of $\mathrm{SO}(3)$, one observes the emergence of an accidentally flat direction parametrized by Eq. (2). The masses of the complex $\chi$ components are given by Eq. (4). One-loop corrections give rise to the usual potential along the accidentally flat direction $V_{\rm eff}=V_{0}^{\prime\prime}-\left(M^{\prime\prime}\right)^{4}\cos\left(a/f\right)$, with $M^{\prime\prime}$ given by
$\left(M^{\prime\prime}\right)^{4}=\frac{1}{160\,\pi^{2}}\left[9\,g^{4}\,v^{4}+4\,\frac{\tilde{\mu}_{\chi}^{10}}{\zeta^{3}\,v^{6}}\left({\rm Re}\;F\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)-T_{F,5}\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)\right)\right]\,,$ (B2)
@@
$\chi_{R}^{a}\left[-\mu_{\chi}^{2}+\lambda_{\chi}\left(\chi_{R}^{2}+\chi_{I}^{2}\right)+(\kappa+\delta)(\chi_{R}^{2}-\chi_{I}^{2})\right]-2\chi_{I}^{a}\left(\chi_{R}\cdot\chi_{I}\right)\left(\kappa-\delta\right)$ $=0\,,$ (B3)
@@
$(\chi_{R}^{3})^{2}=(\chi_{I}^{3})^{2}=w^{2}\equiv\frac{\mu_{\chi}^{2}}{2\left(\lambda_{\chi}+\delta-\kappa\right)}\,,\quad\chi_{I}^{2}=0\,.$ (B5)
In these minima the symmetry is spontaneously broken as $\mathrm{SO}(3)\times\mathbb{Z}_{4\chi}\to\mathrm{SO}(2)\times\mathbb{Z}_{2\chi}$. There are four solutions to Eq. (B5) but only two physically distinct vacua: one (+)(+)( + ) where the $\chi_{R}$ and $\chi_{I}$ VEVs are parallel, $\chi_{R}^{3}=\chi_{I}^{3}=\pm w$, and the other ()(-)( - ) where they are antiparallel, $\chi_{R}^{3}=-\chi_{I}^{3}=\pm w$. Thus, the vacuum manifold has two disconnected components, and DWs can be produced during the SSB phase transition, as discussed in section 5.
For a viable phenomenology, the DW network needs to be unstable, as discussed in section 5.1. We therefore assume that $\mathbb{Z}_{4\chi}$ is softly broken by a term $V_{\rm soft}=im_{\chi}^{2}\chi\chi+{\rm h.c.}$, with $m_{\chi}^{2}$ real. This term preserves $\mathbb{Z}_{2\chi}:\chi\to-\chi$, while it breaks the CP symmetry $\chi\to\chi^{*}$. Consequently, it breaks the degeneracy between the two disconnected vacua ($+$) and (--), introducing a bias in the potential. We assume that soft $\mathbb{Z}_{4\chi}$ breaking is a small effect, $m_{\chi}^{2}\ll\mu_{\chi}^{2}$. Then the potential bias is, to leading order in $m_{\chi}^{2}$,
$\Delta V\equiv V_{(+)}-V_{(-)}=-\frac{2\,m_{\chi}^{2}\mu_{\chi}^{2}}{\lambda_{\chi}+\delta-\kappa}+\mathcal{O}\left(m_{\chi}^{4}\right)\,,$ (B6)

### v2

@@ Appendix C A model with disconnected minima, giving rise to unstable domain walls
Let us first consider the case $\mu_{\phi}^{2}>0$ and $\mu_{\chi}^{2}<0$. The minimum is at $\langle\phi^{2}\rangle=v^{2}\equiv\mu_{\phi}^{2}/\lambda_{\phi}$, $\langle\chi\rangle=0$, where the mass spectrum resembles the one of the minimal model of Sec. 2. After the spontaneous breaking of $\mathrm{SO}(3)$, one observes the emergence of an accidentally flat direction parametrized by Eq. (2). The masses of the complex $\chi$ components are given by Eq. (4). One-loop corrections give rise to the usual potential along the accidentally flat direction $V_{\rm eff}=V_{0}^{\prime\prime}-\left(M^{\prime\prime}\right)^{4}\cos\left(a/f\right)$, with $M^{\prime\prime}$ given by
$\left(M^{\prime\prime}\right)^{4}=\frac{1}{160\,\pi^{2}}\Bigg{[}9\,g^{4}\,v^{4}+4\,\frac{\tilde{\mu}_{\chi}^{10}}{\zeta^{3}\,v^{6}}\Bigg{(}{\rm Re}\;F\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)-T_{F,5}\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)\Bigg{)}\Bigg{]}\,,$ (C2)
@@
$\chi_{R}^{a}\left[-\mu_{\chi}^{2}+\lambda_{\chi}\left(\chi_{R}^{2}+\chi_{I}^{2}\right)+(\kappa+\delta)(\chi_{R}^{2}-\chi_{I}^{2})\right]-2\chi_{I}^{a}\left(\chi_{R}\cdot\chi_{I}\right)\left(\kappa-\delta\right)$ $=0\,,$ (C3)
@@
$(\chi_{R}^{3})^{2}=(\chi_{I}^{3})^{2}=w^{2}\equiv\frac{\mu_{\chi}^{2}}{2\left(\lambda_{\chi}+\delta-\kappa\right)}\,,\quad\chi_{I}^{2}=0\,.$ (C5)
In these minima, the symmetry is spontaneously broken as $\mathrm{SO}(3)\times\mathbb{Z}_{4\chi}\to\mathrm{SO}(2)\times\mathbb{Z}_{2\chi}$. There are four solutions to Eq. (C5) but only two physically distinct vacua: one (+)(+)( + ) where the $\chi_{R}$ and $\chi_{I}$ VEVs are parallel, $\chi_{R}^{3}=\chi_{I}^{3}=\pm w$, and the other ()(-)( - ) where they are antiparallel, $\chi_{R}^{3}=-\chi_{I}^{3}=\pm w$. Thus, the vacuum manifold has two disconnected components, and DWs can be produced during the SSB phase transition, as discussed in Sec. 5.
For a viable phenomenology, the DW network needs to be unstable, as discussed in Sec. 5.1. We, therefore, assume that $\mathbb{Z}_{4\chi}$ is softly broken by a term $V_{\rm soft}=im_{\chi}^{2}\chi\chi+{\rm h.c.}$, with $m_{\chi}^{2}$ real. This term preserves $\mathbb{Z}_{2\chi}:\chi\to-\chi$, while it breaks the CP symmetry $\chi\to\chi^{*}$. Consequently, it breaks the degeneracy between the two disconnected vacua ($+$) and (--), introducing a bias in the potential. We assume that soft $\mathbb{Z}_{4\chi}$ breaking is a small effect, $m_{\chi}^{2}\ll\mu_{\chi}^{2}$. Then the potential bias is, to leading order in $m_{\chi}^{2}$,
$\Delta V\equiv V_{(+)}-V_{(-)}=-\frac{2\,m_{\chi}^{2}\mu_{\chi}^{2}}{\lambda_{\chi}+\delta-\kappa}+\mathcal{O}\left(m_{\chi}^{4}\right)\,,$ (C6)